OpenStudy (anonymous):

Find the equation of the tangent line to the curve y=6sinx at the point (π/6,3). The equation of this tangent line can be written in the form y = mx+b where m=? and b=? How do I find m and b?

OpenStudy (cwrw238):

first find the derivative of the function and plug in x = pi/6 - this will give you the slope m then you can use the genaral equation of straight line: y-y1 = m(x-x10 where (x1,y1) = (pi6/3) and rearrange to give the form y = mx+b

OpenStudy (cwrw238):

* thats y-y1 = m(x-x1)

OpenStudy (anonymous):

Thank you very much. I found the answer! m=6cos(π/6) b≈0.279300953649